Fermi National Laboratory
 
Operation Summary for April 25, 2012 - April 27, 2012
 

Wednesday April 25

The day shift began with Operations monitoring the Linac, Booster, and Main Injector (MI), and with the NuMI, SeaQuest, MTA, and the Meson FTBFT-1015 experiments taking beam.

During the day shift, at 9:45 AM, Operations prepared the Muon Ring, Transport, and the Neutrino Area for access. At 10:06 AM, Operations, as a test, began injecting six batches of protons from MI to Switchyard (SY). Operations had the Neutrino Area and Muon Ring back on by 11:23 AM. At 11:52 AM, Operations turned off Booster and prepared it for an access to investigate a beam stop problem; technicians replaced a micro-switch and were out of access in about an hour. A Booster person began a study at 1:31 PM. MI and SeaQuest personnel began a Septa extraction study at 2 PM. A Muon Ring person began a study at 2:37 PM.

During the evening shift, Operations reestablished normal beam to SeaQuest around 4:30 PM. The Muon Ring study ended at 4:41 PM. Booster studies ended at 4:47 PM. SeaQuest made an access during this shift.

Thursday April 26

The midnight shift began with Operations sending beam to NuMI, SeaQuest, MTA, and T-1015.

During the midnight shift, Operations disabled the MTA beam at 12:21 AM. Operations completed at Booster injection tune at 3:29 AM. Operations began tuning the Booster at 4:34 AM. SeaQuest made an access during this shift.

During the day shift, a Booster person began a study at 9:10 AM. At 10:12 AM, the NuMI beam permit tripped off due to a quadrupole magnet power supply’s (E:QF115F) output drifting high; everything reset. A Muon Ring person began a study at 11:12 AM. Operations established beam to MTA at 12:12 PM. Operations prepared the Muon Ring for a one-hour access at 12:54 PM.

During the evening shift, the Booster study ended at 4:27 PM. The Muon Ring study ended at 4:38 PM. Operations disabled beam to MTA at 6:17 PM. T-1015 made two accesses during this shift.

Friday April 27

The midnight shift began with Operations sending beam to NuMI, SeaQuest, and T-1015.

During the midnight shift, Operations reported having some Booster transition problems at 2:13 AM; Operators made adjustments. The NuMI beam permit tripped off at 7:22 AM due to an MI incomplete RF gap, but there were no RF stations off; everything reset.

Operations Coordinator (OpsCo) Plans for Today and the Weekend

The OpsCo agrees with Booster, MI, and Muon Ring study plans for today and the weekend. Beam to all experiments (except MTA) shuts off at midnight on Monday morning. MI will continue to take beam until 6 AM. Linac, the Neutron Therapy Facility, MTA, and Booster will continue using beam for one or two more weeks. The Fermi Accelerator Complex will be in shutdown for approximately one year.

 

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 04/27/2012    email Fermilab
Security, Privacy, Legal  |  Use of CookiesFermi National Accelerator Laboratory